We wish to create a 90% confidence interval for the Variance given that a sample of 30 has a standard deviation of 3. Find the Lower value for the confidence interval.Round to tenths.
Answer #1

Ansi Given N=30 Std=3 Sample Variance (52)=339 Confidence Interval for variance (or) (n-1)32 sos (1-1)3 df = n1 =21 1-0.90 *

(n-1)3= 2909 = 261 hower limit – (n-1) 37 ou sho Ver Desano) ( 10 = 261 420557 -- 6.133 opper limit= (n=1932 = 261 77-7084 bo